OpenStudy (anonymous):

Can you help me find the circumcenter of this triangle? The vertices of triangle are A(2,2), B(4,10), and C(12,2).

OpenStudy (anonymous):

Do you know how to find midpoint?

OpenStudy (anonymous):

Yes, I have found the midpoint already.

OpenStudy (anonymous):

Ok, of two sides? What are they so we can make some equations

OpenStudy (anonymous):

okay I did the midpoint of A and B which is (3,6)

OpenStudy (anonymous):

Alright we need one more so wanna do the midpoint of BC?

OpenStudy (anonymous):

So it would be 4+12/2=8 and 10+2/2=6 so the midpoint is (8,6) right?

OpenStudy (anonymous):

Yep, now we need to find the slope of AB and BC

OpenStudy (anonymous):

So we would have to do y2-y1 and x2-x1?

OpenStudy (anonymous):

the slope of AB= (8,2)

OpenStudy (anonymous):

and the slope of BC= (-8,8) i think?

OpenStudy (anonymous):

Slope isn't a coordinate. 10-2/4-2 8/2 4 slope of AB is 4

OpenStudy (anonymous):

Oh okay so the slope of BC= -8/8

OpenStudy (anonymous):

Yes so -1. so slope of AB is 4, midpoint of AB is (3,6) we need the slope of the perpendicular bisector of ab so we change the sign, and flip it 4 turns into -1/4 now we put this into point slope form y-6=-1/4(x-3) wanna simplify that so its in slope intercept form?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So do we solve for y?

OpenStudy (anonymous):

Yep isolate y on the left y-6=(-1/4)x+3/4 y=(-1/4)x+27/4 ok keep that equation for a second, do the same thing with BC slope=-1 midpoint=(8,6)

OpenStudy (anonymous):

So y-6=(-1)x+8

OpenStudy (anonymous):

Okay, you lost me there. Before I got here, I made it to the point slope form and then got lost. How did you get the 3/4 in y-6=(-1/4)x+3/4? And How did it change to 27/4? And what am I supposed to do after I solve for y?

OpenStudy (anonymous):

Okay so what do we have to do now?

OpenStudy (anonymous):

Figure out what x and y are, which are the coordinates for the circumcenter. do you know how to solve a system of equations?

OpenStudy (anonymous):

No i mean you have two equations, two variables in each and you have to find out what each variable is